-- PaulSmith - 20 Feb 2007

SDT Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith

Inst. Support: Olga Kuhn TO: Aaron Ceranski

1820: Start biases. Do 5 and then 25. Unable to open due to a similar problem with the brake system as was seen over last weekend (system overheating).

1900: Unable to do evening twilight flats as they work on the telescope problem, but the problem seems to be getting fixed.

1910: Dome opened. Sky clear (probably photometric during the entire night).

1912: Focus field at L95_BV. Seeing appears to be quite bad. Trouble focusing. Roof quite warm from getting rid of ice during the day.

1950: Head to GRB field, do Wave Front (WF) then OB=grb20070125; Seeing 0.9-1.5 but some guiding errors apparent on some frames (nothing too serious). OB repeated. WF done (big focus change). Do OB for 3rd and final time.

2115: WF done at new field. Do OB=fuobj20_2_r; seeing 1.1. Finish the OBs for this target firdt 3 OBs done on 20070219).

2145: WF done at M67. OB=WDs_M67_short_BV

2204: OB=WDs_M67_SCI_B3; seeing/image quality starts out at 1.1". This is a long OB, so we'll see how the image holds together.

2240: WFs done to determine which programs can be attempted.

2302: OB=IC2574_B2; seeing 0.9-1.1".

2330: WFs then OB=Egami_Lockman_VLA_B_1; seeing 1.0-1.6" (3rd-5th exposure of OB shows what looks to be a tracking problem: "box-like PSF); do WF and then to OB=Egami_Lockman_XMM_B_1; seeing has gotten worse - 1st exp comes in at 1.7"; 2nd exp 1.4"; 3rd - 5th exps improved, seeing 1.2".

0055: Move to M81 field and do WF.

0059: OB=KochanekM81V; seeing 0.9-1.4". OB=KochanekM81BV; seeing 1.3-1.6".

0127: WF then OB=LeoIV_B; seeing 1.1", WF then OB=LeoIV_2_B; we thought we had lost the first file of the OB (a 300 s exposure never got to /Repository when it was supposed to). However, this file showed up on disk AFTER the 2nd exposure of the OB. The late fits file is 090838. Seeing 1.0".

0230: WF and then a standard star field in UBVgr filters (PG1323).

0242: Go to IC2574 and do WF, then OB=IC2574_U1; seeing 1.0-1.2". WF then OB=IC2574_U2; seeing 0.9-1.0".

0348: Go to M94 and WF (offset -20' for WF), then OB=M94_NGC_4736_U; seeing 0.6-0.8 (at U!!!). WF then OB=M94_NGC_4736_V; seeing 0.7-1.0". Again, last exposure takes a very long time to find its way into /Repository.

0524: WF at standard star field. Standard field (PG1633) done in UBVr filters. The U OB was redone (x5) to get some counts.

0548: Standard field (PG1323). No WF done before OB=PG1323_BV.

0604: Start sky flats at UBVr. Rotator PA=0 deg.

0630: sky flats with Rotator PA=180 deg. Filters done: U (sky too bright).

0640: Close dome. End night.
Topic revision: r3 - 25 Feb 2007, PaulSmith
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback